OpenStudy (anonymous):

Each side of a square is increasing at a rate of 5 cm/s. How fast is the area increasing when the area is 64cm2?

OpenStudy (anonymous):

\[\large\frac{dA}{dt}=\frac{dA}{ds}\cdot\frac{ds}{dt}\]